# countnullcountdimennullnullnulldimennullnullnulldimennullnullnulldimennullnullnulldimen

### v1

@@ Abstract
When selling many goods with independent valuations, we develop a distributionally robust framework, consisting of a two-player game between seller and nature. The seller has only limited knowledge about the value distribution. The seller selects a revenue-maximizing mechanism, after which nature chooses a revenue-minimizing distribution from all distributions that comply with the limited knowledge. When the seller knows the mean and variance of valuations, bundling is known to be an asymptotically optimal deterministic mechanism, achieving a normalized revenue close to the mean. Moving beyond this variance assumption, we assume knowledge of the mean absolute deviation (MAD), accommodating more dispersion and heavy-tailed valuations with infinite variance. We show for a large range of MAD values that bundling remains optimal, but the seller can only guarantee a revenue strictly smaller than the mean. Another noteworthy finding is indifference to the order of play, as both the max-min and min-max versions of the problem yield identical values. This contrasts with deterministic mechanisms and the separate sale of goods, where the order of play significantly impacts outcomes. We further underscore the universality of the optimal bundling price by demonstrating its efficacy in optimizing not only absolute revenue but also the absolute regret and ratio objective among all bundling prices.
@@ 1 Introduction
We study a single buyer, multi-good Bayesian mechanism design problem in which the seller wants to maximize expected revenue, i.e., the profit obtained from selling the goods. The expectation here is taken with respect known prior distributions for the values that the buyer has for the goods. Over the last decades, an extensive and beautiful theory has been developed for such Bayesian mechanism design problems, dating back to the work of Myerson [37] on auctions.
In the case of selling $m=1$ good with non-negative value distribution $\mathbb{P}$, the optimal truthful mechanism [37, 41, 40] maximizing the revenue is characterized as follows: Set the posted “take-it-or-leave-it” price $p^{*}=\text{argmax}_{p}\,p\cdot\mathbb{P}(X\geq p)$ and sell the good to the buyer if and only if the buyer’s (random) value $X$ for the good exceeds $p^{*}$. Truthfulness refers to the fact that the buyer should have no incentive to lie about the valuation they have for the good. When selling $m>1$ goods, the optimal mechanism may exhibit a much more complex structure [22, 12, 34, 11, 27].
The assumption that we have full knowledge about the prior value distributions of the buyer for the goods is often deemed unrealistic in practice. Adding to this is what is commonly known as Wilson’s doctrine [46, 36]: We should aim at designing mechanisms that only require minimal assumptions about the value distribution of the buyer. This leads to the notion of (distributionally) robust revenue maximization, or robust mechanism design, in which the seller only has partial knowledge about the value distributions of the buyer. This partial knowledge is typically modelled as an ambiguity set that consists of all distributions satisfying the given partial knowledge. A prominent example here is to assume moment knowledge about the unknown distribution, such as the mean and variance. This dates back to the work of Scarf [43]. The idea is now that we have a two-player game where the seller first has to choose a selling mechanism, after which nature gets to choose a distribution from the ambiguity set (satisfying the partial knowledge) for which the seller’s revenue is minimized under the chosen mechanism. Research on robust selling mechanisms and auctions has seen a surge in recent years in the mechanism design literature (see, e.g., Section 1.2). The goal is to either obtain robust counterparts of classical Bayesian results, or to derive tractable robust counterparts of multi-good (and/or multi-buyer) problems for which the Bayesian problem itself has a complex, intractable structure.
Recently there has been an increasing interest in designing robust selling mechanism for the single buyer, multi-good setting [23, 16, 28, 18] (these works are discussed in Section 1.2). In this paper we focus on the setting of robustly selling $m\rightarrow\infty$ independently distributed goods, whose distributions come from a common ambiguity set, to an additive buyer using a deterministic mechanism. This is an important special case of selling multiple goods whose classical Bayesian counterpart, the independent and identical distribution (i.i.d) setting, has received considerable attention in the literature: Bakos and Brynjolfsson [6] show, using the law of large numbers, that if the known distribution is independent of $m$, then bundling [^1] Setting one price for the grand bundle of goods and selling the bundle as a whole if the sum of the valuations of the buyer exceeds the price. is asymptotically (i.e., when $m\rightarrow\infty$) optimal. This can be achieved by setting a price just below $m\cdot\mu$, where $\mu$ is the expectation of the distribution. When the distribution is allowed to depend on $m$, bundling is no longer guaranteed to be optimal [30]. Hart and Nisan [30] showed that, instead, bundling achieves a $c/\log(m)$-fraction of the revenue of the optimal mechanism for some absolute constant $c>0$. Later, Li and Yao [33] improved this result to show that bundling guarantees a constant fraction of the revenue of the (unknown) optimal mechanism.
In the robust setting, if one assumes to know the mean and variance of the unknown distributions, it follows by an application of Chebyshev’s bound that bundling with a price just below $m\mu$ is asymptotically optimal (with a revenue of roughly $m\mu$). This observation can also be seen as a special case of results on robustly selling multiple goods in [18, 28]. These works in particular rely on the assumption of fixing (or upper bounding) the variance of the unknown distributions. In this work we go beyond this assumption and look at a more general ambiguity set that replaces the variance by the mean absolute deviation, as will be explained next. As our main result, we show that (deterministic) bundling is still asymptotically optimal, but in this case it is no longer possible to achieve a revenue close to $m\mu$, creating a stark contrast with the finite-variance case.
@@
The joint distribution $\mathbb{P}^{m}=\Pi_{i=1}^{m}\mathbb{P}_{i}\in\mathcal{P}(\mu,d)^{m}$ is the product of the marginal distributions of all goods where each $\mathbb{P}_{i}$ is from the ambiguity set $\mathcal{P}(\mu,d)$ that assumes to know the mean $\mu$ and mean absolute deviation (MAD) $d$ of the distribution (referred to as mean-MAD ambiguity), i.e.,
@@
Note that while all marginal distributions are contained in the same ambiguity set, they are not assumed to be identical. Hence, the joint distribution $\mathbb{P}^{m}=\Pi_{i=1}^{m}\mathbb{P}_{i}$ may consist of different (marginal) $\mathbb{P}_{i}\in\mathcal{P}(\mu,d)$. The set $\mathcal{P}(\mu,d)$ is non-empty if and only if $0\leq d<2\mu$. The MAD is a natural dispersion measure that can substitute for the variance (or standard deviation) and is less sensitive to outliers than the variance. Furthermore, defining the ambiguity set through the MAD allows distributions with heavy tails that have infinite variance. Ambiguity sets based on the MAD have recently gained more interest in the robust revenue maximization literature, see, e.g., [42, 24, 19].
There is ample empirical evidence for heavy-tailed valuations and demand, also spurred by “the long-tail phenomenon” seen in online retail where niche products often determine a large share of the total sales [21]. Examples of reported heavy tails include demand for books [26], movies [9], spare parts [38], and many more [20], and in the majority of cases the tail exponent is such that the mean is finite while the variance is infinite. The ambiguity set (1.1) we consider in this paper conditions on MAD, and hence allows for such heavy-tailed distributions.
The robust revenue maximization problem we want to solve can be seen as a two-player game, in which the seller (first player) needs to choose a truthful mechanism $A\in\mathcal{A}_{m}$ from the set $\mathcal{A}_{m}$ of allowed truthful mechanisms, after which nature (the second player) chooses a distribution $\mathbb{P}^{m}=\Pi_{i=1}^{m}\mathbb{P}_{i}\in\mathcal{P}^{m}$ with all (independent) marginal distributions $\mathbb{P}_{i}\in\mathcal{P}$, and with the goal of minimizing the expected revenue $\text{REV}(A,\mathbb{P})$ that the seller obtains when using mechanism $A$ to sell $m$ independent goods with random values $X_{i}\sim\mathbb{P}_{i}$ for $i=1,\dots,m$. Because in this work we want to let $m$ grow large, we will in fact maximize the revenue scaled by $m$. That is, we want to solve the maximin problem
$\lim_{m\rightarrow\infty}\sup_{A\in\mathcal{A}_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}^{m}}\,\frac{\text{REV}(A,\mathbb{P}^{m})}{m}.$ (1.2)
@@
$\lim_{m\rightarrow\infty}\inf_{\mathbb{P}^{m}\in\mathcal{P}^{m}}\sup_{A\in\mathcal{A}_{m}}\,\frac{\text{REV}(A,\mathbb{P}^{m})}{m}$ (1.3)
in which the roles of the players are reversed: First nature gets to choose a distribution, after which the seller has to choose an optimal selling mechanism for that distribution.
@@
Roughly speaking, the difference between the two problems lies in the fact that in the maximin problem, the seller needs to hedge against distributional uncertainty, whereas in the minimax problem the seller has to hedge against a (known) worst-case distribution.
1.1 Our contributions
As our main result we show that, when $0<d\leq\mu$, both problems (1.2) and (1.3) have the common value $\mu-d/2$. This is summarized in Theorem 1.1. The mechanism solving the maximization problems is a bundling mechanism with a bundling price just below $m(\mu-d/2)$. It is possible to extract a convergence rate in terms of $m$ from our analysis, but leave this to the interested reader.
Theorem 1.1.
Let $0<d<\mu$, and $\mathcal{D}_{m}$ the set of all deterministic truthful mechanisms for selling $m$ goods. Then
$\lim_{m\rightarrow\infty}\displaystyle\sup_{D\in\mathcal{D}_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\,\frac{\mathrm{REV}(D,\mathbb{P}^{m})}{m}=\lim_{m\rightarrow\infty}\displaystyle\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\sup_{D\in\mathcal{D}_{m}}\,\frac{\mathrm{REV}(D,\mathbb{P}^{m})}{m}=\mu-d/2.$ (1.4)
We next summarize the main insights regarding Theorem 1.1.

### v2

# nullnullnulldimennullnullnulldimen

@@ Abstract
Buyer valuations for a collection of goods may exhibit heavy tails, while sellers often know only limited information about their distribution. In such settings, full distributional assumptions or reliance on second moments may be inappropriate. We assume independence across goods and knowledge of the mean and mean absolute deviation (MAD) of valuations, which keeps the model well defined even under heavy-tailed uncertainty. Within this distributionally robust framework, the seller chooses a selling mechanism to maximize worst-case expected revenue. We show that, in the many-goods limit, no selling mechanism can guarantee extraction of the full mean revenue per good. This contrasts sharply with the finite-variance case, where deterministically selling all goods as a single bundle achieves full mean extraction asymptotically. Despite this shortfall, we show that deterministic bundling remains optimal. Moreover, we show that this bundling price is also optimal when restricting attention to deterministic bundling mechanisms under the competitive ratio and absolute regret objectives, leading to a simple and universal pricing rule.
@@ 1 Introduction
In many markets for digital goods firms increasingly rely on bundling rather than selling goods exclusively on an à-la-carte basis. Platforms that historically emphasized individual purchases now often offer access to large collections of goods through bundled pricing schemes, such as subscription-based access to movies, music, or video games. As a result, understanding bundling as a selling mechanism has become an important topic in economics and mechanism design, and a substantial literature has studied its revenue properties and conditions for optimality. A common feature of these markets is the presence of large catalogs of goods combined with negligible marginal costs of supply. While bundling is already economically interesting with only a few goods, the size of modern catalogs naturally raises the question of how revenue guarantees scale with the number of goods. We therefore study a single-buyer, multi-good revenue maximization problem in the many-goods limit, which allows us to characterize how revenue guarantees scale with the number of goods.
Within this many-goods limiting setting, Bakos and Brynjolfsson (1999) show that when the distributions of the goods are i.i.d., fully known, and do not depend on the number of goods, deterministic bundling becomes optimal. This also means that randomized mechanisms cannot improve the expected revenue in this environment, which they typically can in other settings. Using the law of large numbers together with Chebyshev’s inequality, they establish that the seller’s expected revenue per good converges to the mean, which serves as an upper bound for any truthful selling mechanism. Truthfulness refers to the fact that the buyer should have no incentive to lie about the valuation they have for the good. By suitably truncating the distribution, they further extend this result to heavy-tailed distributions with infinite variance.
The result of Bakos and Brynjolfsson (1999) shows the strength of deterministic bundling when full information on the valuation distributions is available. However, full information is often not available or unreliable and therefore deemed unrealistic in practice. Adding to this is what is commonly known as Wilson’s doctrine (Wilson, 1987; Milgrom, 2004): We should aim at designing mechanisms that only require minimal assumptions about the value distribution of the buyer. This leads to the notion of distributionally robust revenue maximization, in which the seller only has partial knowledge about the value distributions of the buyer, often in the form of summary statistics such as the mean and the variance (see, e.g., Scarf (1958)). We therefore adopt a distributionally robust approach, assuming independent valuations and partial knowledge of the mean and the mean absolute deviation (MAD).
By focusing on the MAD rather than the variance, we depart from much of the existing literature, which typically centers on partial information that ensures finite variance. While this is analytically convenient, this excludes distributions with heavy tails that have infinite variance. There is ample empirical evidence for heavy-tailed valuations and demand, also spurred by “the long-tail phenomenon” seen in online retail, where niche products often determine a large share of the total sales (Das et al., 2021). Examples of reported heavy tails include demand for books (Gaffeo et al., 2008), movies (Bimpikis and Markakis, 2016), spare parts (Natarajan et al., 2018), and many more (Clauset et al., 2009), where the tail behavior frequently yields a finite mean but infinite variance. Such heavy-tailed behavior can substantially influence revenue outcomes when only partial distributional information is known.
Our first contribution shows that, in the many-goods limit, all truthful mechanisms achieve expected revenue per good that is strictly below the mean. This stands in stark contrast to the mean-variance setting (Giannakopoulos et al., 2023), where a concentration argument (Chebyshev’s inequality) shows that deterministic bundling yields the full mean, thereby capturing the full market potential similar to the full-information model. Such arguments break down entirely when the variance may be infinite. We then show, as our second contribution, that deterministic bundling remains optimal within this setting despite not capturing the full market potential. Hence, deterministic bundling remains a powerful mechanism when distributional ambiguity and heavy-tails interact.
We also consider deterministic bundling for the competitive ratio and absolute regret objectives, which compare the expected revenue with the optimal expected revenue under full information. The competitive ratio evaluates this comparison in relative terms, whereas absolute regret evaluates it in absolute terms. As our third contribution, we show that, among all deterministic bundling prices, the same price is optimal for expected revenue, competitive ratio, and absolute regret in the many-goods limit. An important motivation for us to consider these alternative objectives is a recent study by Anunrojwong et al. (2024), which investigates robust decision making across all of the aforementioned criteria for the single-good setting. Their central insight is that mechanisms optimized for one objective often perform poorly under another, revealing a form of objective overfitting. They propose a unified approach that directly optimizes for performance across all objectives simultaneously. Our results contribute to this line of inquiry by offering a first step toward understanding objective alignment in a canonical high-dimensional setting.
From a technical perspective, establishing our results requires addressing optimization problems that, because the goods are independent, are much harder than its correlated counterparts. Under correlated valuations, one can resort to well-known primal-dual semi-infinite linear programming techniques or computing a saddle point of a zero-sum game. Independence, from an optimization point of view, results in non-linear (and non-convex) problems. A fitting analogy is comparing the difference in computational complexity between computing a mixed Nash equilibrium (difficult) and a correlated equilibrium in general finite games (tractable via a linear program). Independence thus introduces a substantive layer of complexity into the analysis.
1.1 Further related work
In this section, we discuss related works, starting with the classical Bayesian setting. We then continue with robust mechanisms for selling multiple goods in other settings, as well as the literature that addresses the case when there are multiple buyers. Finally, we discuss related works for robustly selling a single good. These works demonstrate the growing interest in obtaining distributionally robust solutions to classical mechanism design problems, to which our work contributes.
Mechanisms for selling multiple goods in the full-information setting. In the case of selling $m=1$ good with a given valuation distribution, the optimal mechanism is a posted "take-it-or-leave-it" price (Myerson, 1981; Riley and Samuelson, 1981; Riley and Zeckhauser, 1983) which sells the good if and only if the buyer’s (random) valuation for the good exceeds this price. When selling $m>1$ goods, the optimal mechanism may exhibit a much more complex structure (Daskalakis et al., 2017; Giannakopoulos and Koutsoupias, 2018). Natural choices like bundling all goods together or selling them separately are in general not optimal for any distribution. Furthermore, the optimal mechanism need not be deterministic, as opposed to the single-good case. These observations are nicely summarized by (Hart and Nisan, 2017, Examples 1-4). In the many-goods setting with $m\xrightarrow{}\infty$, Bakos and Brynjolfsson (1999) show that bundling is optimal as long as the valuation distribution does not depend on $m$. When the distribution is allowed to depend on $m$, bundling is no longer guaranteed to be optimal. Instead, Hart and Nisan (2017) show that bundling achieves a $c/\log(m)$-fraction of the revenue of the optimal mechanism for some constant $c>0$. Later, Li and Yao (2013) improved this result to show that bundling guarantees a constant fraction of the revenue of the (unknown) optimal mechanism. For finite $m$, Daskalakis et al. (2017) provide a primal-dual framework that allows them to characterize the optimal mechanism for selling a fixed number of $m\geq 2$ goods, thereby also generalizing earlier works such as Manelli and Vincent (2006); Cai et al. (2012); see also Cai et al. (2019) for a more general framework. Daskalakis et al. (2017) also provide a characterization for when bundling is optimal. As a concrete application of this characterization, they consider i.i.d. goods with uniform distribution on $[c,c+1]$ for some scalar $c$. They show that if $m$ is kept fixed, then for $c$ large enough the bundling mechanism is optimal (this was shown by Pavlov (2011) for $m=2$). However, for $c$ fixed, it is shown that for $m$ large enough, the bundling mechanism is not optimal. In general, there is not a simple economic interpretation of the characterization in Daskalakis et al. (2017). There have also been other works concerned with deriving conditions on the known independently distributed goods for when bundling is the optimal mechanism, such as Haghpanah and Hartline (2021); Pavlov (2011); Daskalakis et al. (2017); Menicucci et al. (2015); Manelli and Vincent (2006); Ibragimov and Walden (2010). See (Che and Zhong, 2021, Footnote 1) for an overview.
Robust mechanisms for selling multiple goods. Carroll (2017) introduces a robust perspective for selling multiple goods. It is shown that when correlations between the goods are allowed, the optimal robust mechanism is to sell every good separately. See also Carroll (2019) for an interesting survey. Che and Zhong (2021) generalize Carroll’s work by introducing a partition structure. They then assume to know the mean of every good, and some dispersion information about every bundle (not every individual good within a bundle) in the partition, but not the possible correlations between the goods in different parts of the partition. Deb and Roesler (2021) consider a robust setting related to ours, but instead of moment information assume to have only support information for the unknown distributions, i.e., they only assume that the distributions have positive probability mass on a given interval. They prove that randomized bundling is the optimal mechanism for selling $m$ goods under an exchangeability assumption (that also applies to independent distributions); see also Che and Zhong (2021) for related results in this direction. Giannakopoulos et al. (2023) consider the robust ratio objective and provide various tight, up to constants, results for selling multiple goods under mean-variance ambiguity for both correlated and independent value distributions. See also Chen et al. (2022) for results in this direction. There are also works that study (versions of) revenue maximization with respect to distance-based ambiguity sets, such as Brustle et al. (2020) who use the Prokhorov and Lévy distances in learning-based settings.
Robust mechanisms for multiple buyers. There are also various papers that study the setting with multiple buyers where there are unknown distributions coming from (moment-based) ambiguity sets. Anunrojwong et al. (2022) study the problem of selling one good to $n$ potential buyers, where the unknown (joint) distribution is only assumed to have bounded support, meaning an upper bound on the valuations is known. Bachrach and Talgam-Cohen (2022) consider a similar model for selling a single good to $n$ i.i.d. (potential) buyers, but in addition assume to know the mean of the common distribution. In both these works (in Bachrach and Talgam-Cohen (2022) only for $n=2$), a robust version of the second price auction with (randomized) reserve price arises as the optimal robust solution, which is proved by using a saddle-point argument for the corresponding robust zero-sum game between the seller and Nature. For other works in the multi-buyer setting, see, e.g., Koçyiğit et al. (2020); Allouah and Besbes (2020); Carroll (2019).
Robust mechanisms for single good. Maximin analysis for mean-variance ambiguity was pioneered in Azar and Micali (2012), generalized to higher moments in Carrasco et al. (2018); see also Chen et al. (2022); Suzdaltsev (2018). Other forms of knowledge were also studied, such as percentile and quantile-based information (Eren and Maglaras, 2010; Ge et al., 2025), mean absolute deviation (Roos et al., 2021; Elmachtoub et al., 2021; Chen et al., 2024), or knowing that the valuation distribution is within the proximity of a given reference distribution (Bergemann and Schlag, 2011; Chen et al., 2024). Instead of maximin expected revenue, alternative objectives studied in the literature include absolute regret (Bergemann and Schlag, 2008, 2011; Elmachtoub et al., 2021; Chen et al., 2024) and the competitive ratio (Eren and Maglaras, 2010; Giannakopoulos et al., 2023; Wang, 2025). We also refer to the work of Allouah et al. (2023) who assume to have prior knowledge about the probability of sale of a good. In some works in the literature, randomized algorithms are also considered and shown to be optimal in certain cases. For selling one good in the robust setting, Chen et al. (2024) show, based on a functional version of Von Neumann’s minimax theorem (Borwein and Zhuang, 1986), that if randomized mechanisms are allowed in the maximin problem, the values of the maximin and minimax problems are equal to each other. We are not aware of a similar result for the problem of selling multiple goods, in particular not with the additional assumption of independent goods. This relation fails to hold if only deterministic mechanisms are allowed.
1.2 Outline
Section 2 introduces the model and gives the necessary background on mechanism design and distributionally robust optimization. Section 3 presents our main results, stated in three theorems. Their proofs are presented in Sections 4, 5, and 6, respectively. We then reflect on the single-good setting in Section 7. Finally, Section 8 offers concluding remarks and discusses possible avenues for future research.
2 Model and preliminaries
We will start by formally defining deterministic truthful mechanisms and the bundling mechanism. We then introduce the distributionally robust framework using probability theory.
2.1 Mechanism design
A direct revelation mechanism $A$ in the setting of a single buyer that bids on $m$ goods is defined by a pair $(z,\pi)$. Here $z:\mathbb{R}_{\geq 0}^{m}\rightarrow[0,1]^{m}$ is the (randomized) allocation rule, and $\pi:\mathbb{R}_{\geq 0}^{m}\rightarrow\mathbb{R}_{\geq 0}$ the payment rule. For a given valuation vector $v=(v_{1},\dots,v_{m})$ with $v_{i}$ the value the buyer has for good $i$, they receive the goods $i$ with probability $z_{i}(v)$, and get charged $\pi(v)$ in total for all the goods they receive. The utility of the buyer under mechanism $A$ with valuation vector $v$ is then given by $u(A,v)=\langle z(v),v\rangle-\pi(v)$ with $\langle x,y\rangle=\sum_{i=1}^{m}x_{i}y_{i}$ the inner product for two vectors $x,y\in\mathbb{R}^{m}$.
A mechanism $A=(z,\pi)$ is truthful if the following conditions are satisfied:
The first condition ensures that the buyer has no incentive to misreport the true values $v$ that they have for the goods; that is, bidding truthfully is a dominant strategy that maximizes their utility. We refer to this property as incentive compatibility. The second condition guarantees that if a buyer truthfully reports their values, then their utility (or surplus) is non-negative, i.e., they don’t lose anything by participating in the selling mechanism. This property is called individual rationality. We restrict to truthful mechanisms throughout the paper and denote the set of all such mechanisms for $m$ goods by $\mathcal{A}_{m}$. Among these mechanisms, deterministic bundling is of particular importance for our analysis.
Deterministic bundling. A deterministic bundling mechanism $D_{\textup{bund}}$ sells either all goods together at a bundling price if the sum of the valuations meets or exceeds this price; otherwise, it sells none of the goods. To be precise, it sets a price $p_{m}=:\pi(v)$ and defines
$z(v)=\left\{\begin{array}[]{ll}(1,1,\dots,1)&\text{ if }\sum_{i}v_{i}\geq p_{m}\\ (0,0,\dots,0)&\text{ if }\sum_{i}v_{i}<p_{m}\end{array}\right.$.
2.2 Distributionally robust framework
In this work, we assume that the valuations of the buyer are random, denoted by the random vector $X=(X_{1},\dots,X_{m})$ over which we have a joint probability distribution $\mathbb{P}^{m}=\Pi_{i=1}^{m}\mathbb{P}_{i}$. We assume that the $X_{i}$ are independently distributed and indicate their marginal distribution by $\mathbb{P}_{i}$, i.e., $\mathbb{P}^{m}$ is the product distribution $\mathbb{P}_{1}\times\mathbb{P}_{2}\times\dots\times\mathbb{P}_{m}$. The (expected) revenue of a mechanism $A\in\mathcal{A}_{m}$ used for selling $m$ independent goods $(X_{1},\dots,X_{m})\sim\mathbb{P}^{m}$ is given by
$\mathrm{REV}(A,\mathbb{P}^{m})=\mathbb{E}_{\mathbb{P}^{m}}[\pi(X_{1},\dots,X_{m})].$ (2.1)
$\text{BUND}(p_{m},\mathbb{P}^{m})=\mathrm{REV}(D_{\textup{bund}},\mathbb{P}^{m})=p_{m}\cdot\mathbb{P}^{m}\left(\sum_{i=1}^{m}X_{i}\geq p_{m}\right).$ (2.2)
@@ 2.2 Distributionally robust framework
The robust revenue maximization problem we want to solve can be seen as a two-player game, in which the seller (first player) needs to choose a truthful mechanism $A\in\mathcal{A}_{m}$ with the aim of maximizing expected revenue of selling $m$ goods, after which Nature (the second player) chooses a distribution $\mathbb{P}^{m}=\Pi_{i=1}^{m}\mathbb{P}_{i}\in\mathcal{P}(\mu,d)^{m}$ with the aim of minimizing (normalized) expected revenue; see (2.8) below. This joint distribution $\mathbb{P}^{m}=\Pi_{i=1}^{m}\mathbb{P}_{i}$ is the product of the marginal distributions of all $m$ goods where each $\mathbb{P}_{i}$ is from the ambiguity set $\mathcal{P}(\mu,d)$ that assumes to know the mean $\mu$ and mean absolute deviation (MAD) $d$ of the distribution (referred to as mean-MAD ambiguity), i.e.,
@@
The ambiguity set $\mathcal{P}(\mu,d)$ is non-empty if and only if $0\leq d<2\mu$. Since our results are monotone in the parameters $\mu$ and $d$, they continue to hold when $\mu$ is replaced by a lower bound and $d$ by an upper bound. Furthermore, although each marginal distribution is contained in the same ambiguity set, we do not assume them to be identical, so the product joint distribution $\mathbb{P}^{m}=\Pi_{i=1}^{m}\mathbb{P}_{i}$ may consist of different (marginal) $\mathbb{P}_{i}\in\mathcal{P}(\mu,d)$. We remark, however, that all results in this paper remain valid under the additional assumption that the marginals are i.i.d.
We often consider the subclass of two-point distributions with mean $\mu$ and mean absolute deviation (MAD) $d$. This class can be fully parameterized by a single parameter $\alpha\in[d/2\mu,1)$ on the left support point. We denote this family (of two-point distributions) as $\mathcal{P}_{2}(\mu,d)\subset\mathcal{P}(\mu,d)$. For $\alpha\in[d/(2\mu),1)$, the distribution of a random variable $X(\alpha)\sim\mathbb{P}_{\alpha}\in\mathcal{P}_{2}(\mu,d)$ is given by
$X(\alpha)=\left\{\begin{array}[]{ll}x(\alpha)=\displaystyle\mu-\frac{d}{2\alpha}&\text{ w.p. }\alpha\\ y(\alpha)=\mu+\displaystyle\frac{d}{2(1-\alpha)}&\text{ w.p. }1-\alpha\end{array}\right.$. (2.6)
When $X_{1},\dots,X_{m}$ are i.i.d. according to $\mathbb{P}_{\alpha}$, then $Y=\sum_{i=1}^{m}X_{i}\sim\mathbb{P}^{m}_{\alpha}$ with
Because in this work we want to let $m\xrightarrow{}\infty$, we will in fact maximize the revenue scaled by $m$. That is, we want to solve the maximin problem
$\sup_{A\in\mathcal{A}_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\,\frac{\text{REV}(A,\mathbb{P}^{m})}{m}.$ (2.8)
@@
$\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\sup_{A\in\mathcal{A}_{m}}\,\frac{\text{REV}(A,\mathbb{P}^{m})}{m}$ (2.9)
@@
in which the roles of the players are reversed: First nature gets to choose a distribution, after which the seller has to choose an optimal selling mechanism for that distribution. Roughly speaking, the difference between the two problems lies in the fact that in the maximin problem, the seller needs to hedge against distributional uncertainty, whereas in the minimax problem the seller has to hedge against a (known) worst-case distribution. While for $m=1$, it is known that the value of (2.8) and (2.9) are equal if the ambiguity set is convex (Chen et al., 2024), this is generally not known for $m>1$. Furthermore, equality is typically not true if one would restrict to deterministic mechanisms (we return to these observations later).
3 Main results
We first establish an upper bound of $\mu-d/2$ on achievable (normalized) revenue in the maximin problem (2.8) as $m\rightarrow\infty$, showing that no truthful mechanism can extract the full market potential $\mu$ asymptotically. We do this by showing the stronger result that the minimax problem in (2.9), that upper bounds the maximin problem, is upper bounded by $\mu-d/2$.
We then analyze the optimal revenue achievable by robust deterministic bundling and show that it also equals $\mu-d/2$ as $m\rightarrow\infty$. Because this value is a lower bound on the maximin problem, this yields that the maximin and minimax problem are asymptotically equal to each other with a common value of $\mu-d/2$. We conclude by examining the power of deterministic bundling by showing that the optimal bundling price for the expected revenue objective remains optimal among all deterministic bundling mechanisms for the competitive ratio and absolute regret objective.
3.1 Selling below the mean